# Model Informed Flows for Bayesian Inference of Probabilistic Programs

### v1

@@ 3.3 The Necessity of Conditioning on Previously Generated Latent Variables
Inverse Autoregressive Flows (IAFs) (kingma2016improved, ) represent an important class of normalizing flows, lauded for the computational efficiency of their sampling process. Unlike FAFs, conditioning networks of IAFs depend solely on the base noise variables $\epsilon_{1:i-1}$, allowing parallel computation of all $z_{i}$:
@@
$m_{i}(\epsilon_{1:i-1})$ =?${{=}}f_{i}(\pi(z_{i}))+g_{i}(\pi(z_{i}))^{1-\lambda_{i}}\Bigl{(}\mu_{i}+\sum_{j=1}^{i-1}L_{ij}\epsilon_{j}-\lambda_{i}f_{i}(\pi(z_{i}))\Bigr{)},$ (3.13)
$\log s_{i}(\epsilon_{1:i-1})$ =?${{=}}(1-\lambda_{i})\log g_{i}(\pi(z_{i}))+\log L_{ii},$ (3.14)
@@
Particularly, an IAF with affine conditioning networks for $m_{i},\log s_{i}$ (and a term like $t_{i}$, if added) cannot capture the VIP transformation, since $z_{1:i-1}$ and $\epsilon_{1:i-1}$ have a complex non-linear relationship. This limitation underscores the advantage of forward autoregressive structures. We show in our experiments (e.g., results related to Figure 2) that IAFs require substantially greater model capacity to approach the effectiveness of MIF in representing VIP.
Table 1: Negative ELBOs ($-\mathrm{ELBO}$) for hierarchical benchmark models using mean-field Gaussian (MF), mean-field Gaussian with VIP (MF-VIP), full-rank Gaussian (FR), and full-rank Gaussian with VIP (FR-VIP). Lower values indicate tighter posterior approximations.
@@ 4.1 Experimental Setup
For all experiments, we utilize the Adam optimizer (kingma2014adam, ) and initialize all learnable parameters from a standard Gaussian distribution with a standard deviation of 0.1. For our Model-Informed Flow (MIF) evaluations, distinct configurations are used: the experiments in Section 4.3 primarily consider a single-layer affine MIF to clearly assess its structural components. In subsequent experiments, including the network expressiveness analysis and comprehensive benchmarks, MIF’s conditioning networks are implemented as multi-layer perceptrons with ReLU activation functions to explore the impact of increased representational capacity via hidden units.
We perform 100,000 optimization iterations, using 256 Monte Carlo samples to approximate the ELBO during training. After training, the final ELBO is evaluated using 100,000 fresh samples for reliable estimation. For each model configuration, we report the best ELBO achieved after exploring six learning rates (logarithmically spaced from $10^{-1}$ to $10^{-6}$). Additional details regarding the benchmark models, specific variational family configurations, and MLP architectures are provided in Appendix D.
@@ 4.2 Preliminary Observations: Full-Rank VIP Outperforms Baselines
Prior work on VIP focused exclusively on mean-field Gaussian approximations (gorinova2020automatic, ). To assess whether the advantages of VIP extend to richer variational families, we apply VIP to a full-rank Gaussian setting and compare four configurations—MF, MF-VIP, FR, and FR-VIP—across six canonical hierarchical models. As shown in Table 1, FR-VIP achieves the lowest negative ELBO in every case, outperforming both its non-VIP counterpart and mean-field VIP, and achieving performance comparable to more complex state-of-the-art methods (cf. Table 3). These results highlight the power of combining full-rank with VIP and motivate our design of VIP-inspired low architectures.
@@ 4.3 Impact of Design Components
In this section, we empirically validate the contribution of MIF’s distinct design elements through a targeted ablation study, with results presented in Table 2. We compare the affine MIF against variants that individually alter or remove its key theoretically-motivated components: the $\epsilon$-dependent translation term ($t_{i}$), by evaluating MIF (w/o $t_{i}$); the use of explicit prior information, by assessing MIF (w/o Prior) which omits $f_{i},\log g_{i}$ as inputs; and adherence to the hierarchical variable processing order, examined via MIF (w/o Order). Additionally, we test an IAF-like variant, MIF ($\epsilon$-cond), which conditions primarily on base noise $\epsilon_{1:i-1}$ instead of previously generated latents $z_{1:i-1}$, alongside relevant baselines like FR-VIP and a standard IAF. The consistent outcome is that deviating from the full MIF design by removing or altering these components generally degrades performance, underscoring their collective importance to its effectiveness.
Table 2: Negative ELBOs ($-\mathrm{ELBO}$) for ablation study comparing affine Model-Informed Flow (MIF) against variations and baselines across hierarchical models. MIF includes z-conditioning, translation term, prior info, and correct order. Please refer to Appendix D for more details of variants of MIF. Lower values indicate tighter posterior approximations.
Model FR-VIP MIF MIF ($\epsilon$-cond) MIF (w/o $t_{i}$) MIF (w/o Prior) MIF (w/o Order) IAF
@@ 4.4 Impact of Network Expressiveness
Recall from Section 3.3 that while IAFs face fundamental challenges in representing the VIP mechanism due to their conditioning solely on noise ($\epsilon$), it was posited that sufficient network expressiveness might partially overcome these limitations. We empirically investigate this by comparing our full MIF, which conditions on previously generated latents $z_{1:i-1}$, against MIF ($\epsilon$-cond), an IAF-like variant where the conditioning networks primarily use $\epsilon_{1:i-1}$ instead of $z_{1:i-1}$.
Figure 2 illustrates the performance of MIF versus MIF ($\epsilon$-cond) as the expressiveness of their conditioning networks (e.g., by increasing the number of hidden units) is varied. Across most benchmarked models, enhancing network capacity allows MIF ($\epsilon$-cond) to substantially improve its ability to capture complex dependencies. Consequently, the performance gap between this IAF-like variant and the standard MIF significantly narrows with increased expressiveness, demonstrating that sufficiently powerful IAF-like architectures can indeed achieve strong results, in many cases closely approaching or nearly matching the performance of MIF, which itself may maintain a slight lead or offer benefits in parameter efficiency.
Figure 2: Best ELBO achieved by full MIF (blue) and the variant without latent-variable conditioning (orange) as a function of MLP hidden-unit count. Higher capacity allows the no-latent variant to close the gap, showing that sufficiently expressive networks can implicitly learn dependencies otherwise provided by explicit latent inputs.
@@ 4.5 Comprehensive Evaluation on Hierarchical and Non-hierarchical Benchmarks
To assess its generality and competitive strength, we benchmark MIF with 1024 hidden units per conditioning network, denoted MIF$(h=2^{10})$, against several leading variational inference methods. These baselines, primarily drawn from the recent benchmark by blessing2024beyond and detailed with their acronyms in the caption of Table 3, provide a robust comparison. The evaluation spans the six previously discussed hierarchical models and three additional datasets (Seeds, Sonar, Ionosphere) to test MIF’s broad applicability. Results in the Table 3 show that MIF$(h=2^{10})$ achieves the best or equal-best performance on most of the benchmarks. These strong results, obtained using just a single layer of our expressive MIF transformation, demonstrate that its theory-driven design effectively delivers state-of-the-art performance with notable architectural simplicity.
Table 3: Negative ELBOs ($-\mathrm{ELBO}$) for various VI methods on hierarchical and non-hierarchical benchmarks. Compared methods include: Gaussian Mean-Field (MF) (bishop2006pattern, ), Gaussian Mixture Model VI (GMMVI) (arenz2022unified, ), Sequential Monte Carlo (SMC) (del2006sequential, ), Annealed Flow Transport (AFT) (arbel2021annealed, ), Flow Annealed IS Bootstrap (FAB) (midgley2022flow, ), Continual Repeated AFT (CRAFT) (matthews2022continual, ), and Uncorrected Hamiltonian Annealing (UHA) (thin2021monte, ). Lower values indicate tighter posterior approximations; denotes results taken from blessing2024beyond .
Model MF GMMVI SMC AFT FAB CRAFT UHA MIF$(h=2^{10})$
@@
Related Work.
Early flow-based generative models—including MADE (germain2015made, ), Real NVP (dinh2017density, ), Glow (kingma2018glow, ), and Masked Autoregressive Flow (papamakarios2017masked, )—showed that autoregressive masking, coupling transforms, and invertible $1\times 1$ convolutions yield tractable likelihoods and strong density estimation in high dimensions. Normalizing flows were soon adopted for variational inference, first as post-hoc enrichments of diagonal Gaussians (rezende2015variational, ), then as architectures such as Inverse Autoregressive Flow that scale elegantly to large latent spaces (kingma2016improved, ). To mitigate the “funnel” geometry in hierarchical posteriors, Variationally Inferred Parameters (VIP) introduced gradient-based partial non-centering (gorinova2020automatic, ). Building on VIP, subsequent hybrids such as Automatic Structured Variational Inference (ASVI) (ambrogioni2021automaticstructure, ), Cascading Flows (ambrogioni2021automaticvariational, ), and Embedded-Model Flows (EMF) (silvestri2021embedded, ) combine model-specific structure with flow flexibility. Our Model-Informed Flow continues this line by proving that full-rank VIP can be represented in a forward autoregressive flow augmented with translation and prior inputs, yielding a compact yet expressive variational family for hierarchical models.
@@ VI Variational Inference
$\tilde{z}$ Auxiliary base variable in VIP/NCP, $z=T(\tilde{z})$
@@
$u_{i}$ Input to conditioning networks in MIF: $[z_{1:i-1},f_{i}(\pi(z_{i})),g_{i}(\pi(z_{i}))]$
@@ Appendix B Algorithm
1:Parameters $w$ of the base distribution $q_{w}(\tilde{z})$; VIP parameters $\lambda_{1},\dots,\lambda_{D}$; model prior functions $f_{i}(\cdot)$, $g_{i}(\cdot)$, and parent dependencies $\pi(\cdot)$; dimensionality $D$; a topological ordering for $z_{1},\dots,z_{D}$.
2:Sample $\tilde{z}\sim q_{w}(\tilde{z})$
3:$\log J\leftarrow 0$
@@ Appendix C Proof
Now, the VIP transformation $T_{\lambda}$ is defined in Definition 1 as
@@ D.1 Models
The Eight Schools model [39] estimates treatment effects $\theta_{i}$ for $N_{\text{schools}}=8$ schools, given observed effects $y_{i}$ and known standard errors $\sigma_{i}$.
@@
A logistic regression model [17] with $N_{\text{features}}=K$ predictors $x_{nk}$ for $N_{\text{obs}}$ individuals, predicting binary outcomes $y_{n}$. It uses hierarchical priors on coefficient scales $\tau_{k}$.
@@
Neal’s Funnel distribution [33] in $D=10$ dimensions for $\mathbf{x}=(x_{1},\dots,x_{D})$.
@@
All variables $x_{1},\dots,x_{D}$ are latent. (The variance for $x_{1}$ is $93^{2}=$ 9; the variance for $x_{k},k\geq 2$ is $\exp(x_{1})$.)
@@
A hierarchical linear regression [17] modeling log-radon levels $\log r_{j}$ in $N_{\text{homes}}$ homes. Data includes $x_{j}$ (floor indicator for home $j$), $u_{k}$ (uranium reading for county $k$), and $c_{j}$ (county for home $j$). Priors are inferred from the provided code structure.
@@
The Movielens model [21] is a hierarchical logistic regression for predicting movie ratings. Each rating $n$ is given by a user $u_{n}\in\{1,\dots,M\}$ for a movie described by a binary attribute vector $\mathbf{x}_{n}\in\{0,1\}^{D}$ (with $D=18$ attributes). The rating is $y_{n}\in\{0,1\}$.
@@
The Item Response Theory (IRT) two-parameter logistic (2PL) model [30] is used to model student responses to test items. Let $N_{S}$ be the number of students, $N_{Q}$ be the number of questions (items), and $N_{R}$ be the total number of responses. For each response $r$, $s_{r}$ is the student ID and $q_{r}$ is the question ID.
@@
The Seeds model [13] is a random effects logistic regression for analyzing seed germination data. It models the number of germinated seeds $r_{i}$ out of $N_{i}$ seeds in $G=21$ experimental groups, using group-specific predictors $x_{1,i}$ and $x_{2,i}$.
@@
The Sonar model [20] is a Bayesian logistic regression. It uses $D=61$ features/coefficients $\mathbf{x}$ to classify $N=208$ sonar returns $u_{i}$ into two categories $y_{i}\in\{0,1\}$ (e.g., mines vs. rocks).
@@
The Ionosphere model [40] is a Bayesian logistic regression. It uses $D=35$ features/coefficients $\mathbf{x}$ to classify $N=351$ radar returns $u_{i}$ from the ionosphere into two categories $y_{i}\in\{0,1\}$.

### v2

@@ 3.3 The Necessity of Conditioning on Previously Generated Latent Variables
Inverse Autoregressive Flows (IAFs) (kingma2016improved) represent an important class of normalizing flows, lauded for the computational efficiency of their sampling process. Unlike FAFs, conditioning networks of IAFs depend solely on the base noise variables $\epsilon_{1:i-1}$, allowing parallel computation of all $z_{i}$:
@@
$m_{i}(\epsilon_{1:i-1})$ $\stackrel{{\scriptstyle?}}{{=}}f_{i}(\pi(z_{i}))+g_{i}(\pi(z_{i}))^{1-\lambda_{i}}\Bigl(\mu_{i}+\sum_{j=1}^{i-1}L_{ij}\epsilon_{j}-\lambda_{i}f_{i}(\pi(z_{i}))\Bigr),$ (3.13)
$\log s_{i}(\epsilon_{1:i-1})$ $\stackrel{{\scriptstyle?}}{{=}}(1-\lambda_{i})\log g_{i}(\pi(z_{i}))+\log L_{ii},$ (3.14)
@@
Particularly, an IAF with affine conditioning networks for $m_{i},\log s_{i}$ (and a term like $t_{i}$, if added) cannot capture the VIP transformation, since $z_{1:i-1}$ and $\epsilon_{1:i-1}$ have a complex non-linear relationship. This limitation underscores the advantage of forward autoregressive structures. We show in our experiments (e.g., results related to Figure˜2) that IAFs require substantially greater model capacity to approach the effectiveness of MIF in representing VIP.
Table 1: Negative ELBOs ($\mathrm{ELBO}$) for hierarchical benchmark models using mean-field Gaussian (MF), mean-field Gaussian with VIP (MF-VIP), full-rank Gaussian (FR), and full-rank Gaussian with VIP (FR-VIP). Lower values indicate tighter posterior approximations.
@@
4 Related Work
Early flow-based generative models—including MADE (germain2015made), Real NVP (dinh2017density), Glow (kingma2018glow), and Masked Autoregressive Flow (papamakarios2017masked)—established that autoregressive masking, coupling transforms, and invertible $1{\times}1$ convolutions enable tractable likelihoods with strong high-dimensional density estimation. Normalizing flows were subsequently adapted for variational inference, first as post-hoc enrichments of diagonal Gaussians (rezende2015variational) and later as architectures that scale in latent dimensionality, most notably Inverse Autoregressive Flow (IAF) (kingma2016improved). These developments demonstrated that flexible, invertible parameterizations can substantially tighten the ELBO without sacrificing computational efficiency.
For hierarchical Bayesian models, Variationally Inferred Parameters (VIP) introduced gradient-based partial non-centering to alleviate the funnel geometry typical of deep hierarchies (gorinova2020automatic). Building on VIP’s idea of injecting model structure into the variational family, hybrids such as Automatic Structured Variational Inference (ASVI) (ambrogioni2021automaticstructure), Cascading Flows (ambrogioni2021automaticvariational), and Embedded-Model Flows (EMF) (silvestri2021embedded) couple model-aware parameterizations with flow flexibility. Our Model-Informed Flow continues this line by proving that full-rank VIP can be represented in a forward autoregressive flow augmented with translation and prior inputs, yielding a compact yet expressive variational family for hierarchical models.
@@ 5.1 Experimental Setup
For all experiments, we utilize the Adam optimizer (kingma2014adam) and initialize all learnable parameters from a standard Gaussian distribution with a standard deviation of 0.1. For our Model-Informed Flow (MIF) evaluations, distinct configurations are used: the experiments in Section˜5.3 primarily consider a single-layer affine MIF to clearly assess its structural components. In subsequent experiments, including the network expressiveness analysis and comprehensive benchmarks, MIF’s conditioning networks are implemented as multi-layer perceptrons with ReLU activation functions to explore the impact of increased representational capacity via hidden units.
We perform 100,000 optimization iterations, using 256 Monte Carlo samples to approximate the ELBO during training. After training, the final ELBO is evaluated using 100,000 fresh samples for reliable estimation. For each model configuration, we report the best ELBO achieved after exploring six learning rates (logarithmically spaced from $10^{-1}$ to $10^{-6}$). Additional details regarding the benchmark models, specific variational family configurations, training time comparison, and MLP architectures are provided in Appendix˜D. Our implementation is available at https://github.com/joohwanko/Model-Informed-Flow
@@ 5.2 Preliminary Observations: Full-Rank VIP Outperforms Baselines
Prior work on VIP focused exclusively on mean-field Gaussian approximations (gorinova2020automatic). To assess whether the advantages of VIP extend to richer variational families, we apply VIP to a full-rank Gaussian setting and compare four configurations—MF, MF-VIP, FR, and FR-VIP—across six canonical hierarchical models. As shown in Table 1, FR-VIP achieves the lowest negative ELBO in every case, outperforming both its non-VIP counterpart and mean-field VIP, and achieving performance comparable to more complex state-of-the-art methods (cf. Table 3). These results highlight the power of combining full-rank with VIP and motivate our design of VIP-inspired low architectures.
@@ 5.3 Impact of Design Components
In this section, we empirically validate the contribution of MIF’s distinct design elements through a targeted ablation study, with results presented in Table˜2. We compare the affine MIF against variants that individually alter or remove its key theoretically-motivated components: the $\epsilon$-dependent translation term ($t_{i}$), by evaluating MIF (w/o $t_{i}$); the use of explicit prior information, by assessing MIF (w/o Prior) which omits $f_{i},\log g_{i}$ as inputs; and adherence to the hierarchical variable processing order, examined via MIF (w/o Order). Additionally, we test an IAF-like variant, MIF ($\epsilon$-cond), which conditions primarily on base noise $\epsilon_{1:i-1}$ instead of previously generated latents $z_{1:i-1}$, alongside relevant baselines like FR-VIP and a standard IAF. The consistent outcome is that deviating from the full MIF design by removing or altering these components generally degrades performance, underscoring their collective importance to its effectiveness.
Table 2: Negative ELBOs ($\mathrm{ELBO}$) for ablation study comparing affine Model-Informed Flow (MIF) against variations and baselines across hierarchical models. MIF includes z-conditioning, translation term, prior info, and correct order. Please refer to Appendix˜D for more details of variants of MIF. Lower values indicate tighter posterior approximations.
Model FR-VIP MIF
MIF
($\epsilon$-cond)
MIF
(w/o $t_{i}$)
MIF
(w/o Prior)
MIF
(w/o Order)
IAF
@@ 5.4 Impact of Network Expressiveness
Recall from Section˜3.3 that while IAFs face fundamental challenges in representing the VIP mechanism due to their conditioning solely on noise ($\epsilon$), it was posited that sufficient network expressiveness might partially overcome these limitations. We empirically investigate this by comparing our full MIF, which conditions on previously generated latents $z_{1:i-1}$, against MIF ($\epsilon$-cond), an IAF-like variant where the conditioning networks primarily use $\epsilon_{1:i-1}$ instead of $z_{1:i-1}$.
Figure˜2 illustrates the performance of MIF versus MIF ($\epsilon$-cond) as the expressiveness of their conditioning networks (e.g., by increasing the number of hidden units) is varied. Across most benchmarked models, enhancing network capacity allows MIF ($\epsilon$-cond) to substantially improve its ability to capture complex dependencies. Consequently, the performance gap between this IAF-like variant and the standard MIF significantly narrows with increased expressiveness, demonstrating that sufficiently powerful IAF-like architectures can indeed achieve strong results, in many cases closely approaching or nearly matching the performance of MIF, which itself may maintain a slight lead or offer benefits in parameter efficiency.
Figure 2: Best ELBO achieved by full MIF (blue) and the variant without latent‐variable conditioning (orange) as a function of MLP hidden‐unit count. Higher capacity allows the no‐latent variant to close the gap, showing that sufficiently expressive networks can implicitly learn dependencies otherwise provided by explicit latent inputs.
@@ 5.5 Comprehensive Evaluation on Hierarchical and Non-hierarchical Benchmarks
[…]
Table 3: Negative ELBOs ($\mathrm{ELBO}$) for various VI methods on hierarchical and non-hierarchical benchmarks. Compared methods include: Gaussian Mean-Field (MF) (bishop2006pattern), Gaussian Mixture Model VI (GMMVI) (arenz2022unified), Sequential Monte Carlo (SMC) (del2006sequential), Annealed Flow Transport (AFT) (arbel2021annealed), Flow Annealed IS Bootstrap (FAB) (midgley2022flow), Continual Repeated AFT (CRAFT) (matthews2022continual), and Uncorrected Hamiltonian Annealing (UHA) (thin2021monte). Lower values indicate tighter posterior approximations; denotes results taken from blessing2024beyond.
[…]
@@ 6 Discussion and Limitation
7 Acknowledgement
This material is based upon work supported in part by the National Science Foundation under Grant No. 2045900.
@@ VI Variational Inference
$\tilde{z}$ Auxiliary base variable in VIP/NCP, z=T(z~)z=T($\tilde${z})
@@
$u_{i}$ Input to conditioning networks in MIF: $z_{1:i-1},f_{i}(\pi(z_{i})),g_{i}(\pi(z_{i}))]$
@@ Appendix B Algorithm
[…]
@@ Appendix C Proof
Now, the VIP transformation $T_{\lambda}$ is defined in Definition˜1 as
@@ D.1 Models
The Eight Schools model [rubin1981estimation] estimates treatment effects $\theta_{i}$ for $N_{\text{schools}}=8$ schools, given observed effects $y_{i}$ and known standard errors $\sigma_{i}$.
@@
A logistic regression model [gelman2007data] with $N_{\text{features}}=K$ predictors $x_{nk}$ for $N_{\text{obs}}$ individuals, predicting binary outcomes $y_{n}$. It uses hierarchical priors on coefficient scales $\tau_{k}$.
@@
Neal’s Funnel distribution [neal2003slice] in $D=10$ dimensions for $\mathbf{x}=(x_{1},\dots,x_{D})$.
@@
All variables $x_{1},\dots,x_{D}$ are latent. (The variance for $x_{1}$ is $3^{2}=9$; the variance for $x_{k},k\geq 2$ is $\exp(x_{1})$.)
@@
A hierarchical linear regression [gelman2007data] modeling log-radon levels $\log r_{j}$ in $N_{\text{homes}}$ homes. Data includes $x_{j}$ (floor indicator for home $j$), $u_{k}$ (uranium reading for county $k$), and $c_{j}$ (county for home $j$). Priors are inferred from the provided code structure.
@@
The Movielens model [harper2015movielens] is a hierarchical logistic regression for predicting movie ratings. Each rating $n$ is given by a user $u_{n}\in\{1,\dots,M\}$ for a movie described by a binary attribute vector $\mathbf{x}_{n}\in\{0,1\}^{D}$ (with $D=18$ attributes). The rating is $y_{n}\in\{0,1\}$.
@@
The Item Response Theory (IRT) two-parameter logistic (2PL) model [lord2008statistical] is used to model student responses to test items. Let $N_{S}$ be the number of students, $N_{Q}$ be the number of questions (items), and $N_{R}$ be the total number of responses. For each response $r$, $s_{r}$ is the student ID and $q_{r}$ is the question ID.
@@
The Seeds model [crowder1978beta] is a random effects logistic regression for analyzing seed germination data. It models the number of germinated seeds $r_{i}$ out of $N_{i}$ seeds in $G=21$ experimental groups, using group-specific predictors $x_{1,i}$ and $x_{2,i}$.
@@
The Sonar model [gorman1988analysis] is a Bayesian logistic regression. It uses $D=61$ features/coefficients $\mathbf{x}$ to classify $N=208$ sonar returns $u_{i}$ into two categories $y_{i}\in\{0,1\}$ (e.g., mines vs. rocks).
@@
The Ionosphere model [sigillito1989classification] is a Bayesian logistic regression. It uses $D=35$ features/coefficients $\mathbf{x}$ to classify $N=351$ radar returns $u_{i}$ from the ionosphere into two categories $y_{i}\in\{0,1\}$.